Use the following bond quotes to answer Questions 3-5. To calculate the number of years until maturity, assume that it is currently January 15, 2006. All bonds make semiannual interest payments

company Last last Vol
(ticker) coupon Maturity price Yield (000s)
xenon 6.80 1/15/17 816.584 ?? 64,381
Kenny 8.40 1/15/26 ?? 7.16 43,289
williams ?? 1/15/18 1086.452 8.41 37,653

3. What is the yield to maturity for the bond issued by Xenon?

Can you please explain how to solve the problem on a financial calculator.
Answer #1

Using financial calculator
Input: FV= 1000

PMT = 6.8%*1000/2 = 34

N = 11*2 = 22

PV = -816.584

Solve for I/Y as 4.49

Hence YTM = 4.49* 2

= 8.98%